Bogard,by himself, can paint four rooms in ten hours. If bogard hires April to assist, then they can do the same job together in six hours. If Bogard lets April work alone, how long will it take her to paint the four rooms? (Has to be answered in Algebra form please?)
Answer by stanbon(75887)

Bogard,by himself, can paint four rooms in ten hours. If bogard hires April to assist, then they can do the same job together in six hours. If Bogard lets April work alone, how long will it take her to paint the four rooms?
-----------------
Bogard DATA:
Time = 10 hrs/job ; Rate = 1/10 job/hr.
-----------------------------
Together DATA:
Time = 6 hrs/job ; Rate = 1/6 job/hr
-------------------------------
April alone DATA:
Time = x hrs/job ; Rate = 1/x job/hr
--------------------------------
EQUATION:
Bogard rate + April rate = together rate
1/10 + 1/x = 1/6
3x + 30 = 5x
2x = 30
x = 15 hrs. (time April needs to do the job alone)
=====================================
Cheers,
Stan H.